OpenStudy (anonymous):

If ƒ(x ) = 2x, then ƒ -1(x ) =s

OpenStudy (anonymous):

what do you need to find?

OpenStudy (solomonzelman):

If you have a function \(\Large\color{blue}{ f(x)=\color{orangered}{\rm k}x}\) (where K is a constant) The steps you take in general: 1) replace y with f(x) 2) replace x and y with each other (write x instead of y and vice versa) 3) solve for y (not by replacing it back with x, but by math manipulations) 4) once you isolated y, denote the y as \({\rm f^{-1}(x)}\) ---------------------------------------------------------------- In your case, you are dealing with: \(\large\color{blue}{ f(x)=\color{orangered}{\rm k}x}\) step 1: \(\large\color{blue}{ y=\color{orangered}{\rm k}x}\) step 2: \(\large\color{blue}{ x=\color{orangered}{\rm k}y}\) step 3: \(\large\color{blue}{ \displaystyle\frac{x}{\color{orangered}{\rm k}}=y}\) step 4: \(\large\color{blue}{ \displaystyle\frac{x}{\color{orangered}{\rm k}}={\rm f}^{-1}(x)}\)